# Machine-learning for photoplethysmography analysis: Benchmarking feature, image, and signal-based approaches

### v1

@@ 4.2 Atrial Fibrillation Detection
T XResNet1d101 0.86 0.70 0.76 0.73 0.55 0.52
XResNet1d50 0.87 0.69 0.78 0.78 0.57 0.57
Inception1d 0.87 0.72 0.79 0.79 0.58 0.58
LeNet1d 0.76 0.55 0.58 0.50 0.37 0.32
Alexnet1d 0.84 0.67 0.73 0.71 0.52 0.50
Minirocket 0.82 0.65 0.68 0.67 0.47 0.47
TCN+MLP 0.86 0.68 0.74 0.76 0.54 0.54
F WAVELET + MLP 0.77 0.61 0.59 0.52 0.38 0.33
@@
I CWT 0.82 0.69 0.72 0.69 0.50 0.49
Table LABEL:table_5 represents the performance analysis of the classification task (AF/ non-AF) based on the Deepbeat dataset. Similarly to the regression task, raw time series (T) and image-based (I) models performed better than feature-based (F) models in all cases except for the LeNet1d raw time series model, which performed poorly. The best-performing models were Inception1d and XResNet1d50, which achieved AUCs of 0.87, and F1 scores of 0.72 and 0.69 respectively. As with the regression task, whilst best performance was achieved with the more complex raw time series models, the image-based model performed better than some raw time series models. Whilst feature-based models generally performed worst, wavelet-based features produced better performance than clinically interpretable features.
@@ 5.1 Blood Pressure Estimation
Relative Performance Comparison. For both regression tasks, the best results were achieved by approaches based on raw time series. Feature-based approaches were not competitive in the Calib scenario but are to a certain degree competitive in the CalibFree scenario. Here, it is important to note the differences between Calib and CalibFree: CalibFree tests on unseen patients whereas Calib purposely tests on patients already seen during training, i.e., models can profit to a certain degree from memorization. Quite naturally models achieved much better scores in the Calib setting (even though test sets are not entirely comparable). This also impacts the best-forming models in each of the two settings (within the category of models operating on raw time series). Large and complex models (such as XResNet1d and Inception1d) performed best on Calib, as overfitting to specific patients is desirable, whereas smaller models (such as LeNet1d) performed on par or even better than more complex models in the CalibFree scenario, where generalization to unseen patients is key. This observation aligns with the fact that the gap between raw time series and feature-based approaches is larger in the case of Calib, which is quite natural as this task largely profits from the complexity of the model (and all feature-based approaches are less complex than typical models operating on raw time series).
@@
Comparison to Literature Results. The number of published results on PulseDB is very limited. To the best of our knowledge [46] is the only prior study that reported results on PulseDB, albeit on the full dataset and not just the VitalDB subset, achieving a MAE of 10-11 mmHg for SBP after self-supervised pretraining. We refer to [47] for a recent comparative analysis of different models on different datasets, where models achieved MAEs of between 11 and 19 mmHg for SBP, and between 7 and 11 mmHg for DBP. In comparison, in this study the lowest MAEs were 12.34 and 7.84 mmHg for SBP and DBP respectively on VitalDB CalibFree. This suggests that the presented best-performing models reach state-of-the-art performance.
@@
Impact of Loss Function and Dropout Ensembling. The performance of the XResNet1d50+GNLL in comparison to the XResNet1d50 trained with a regular MAE loss highlights the benefits incorporating dropout ensembling at evaluation, and using a likelihood-based loss can have on model performance. It suggests that potentially further CNN results could be improved by using a different loss function and ensemble-based evaluation procedure. Interestingly, this only applied to the Calib and not the CalibFree scenario. Other works have reported that modelling uncertainties may improve predictive performance on models trained on ECG data [48]. This variant had two parallel output branches with dropout layers (with a constant but low dropout rate, e.g. 4-5% depending on the task/parameter) dispersed throughout the architecture. Each test input was evaluated 100 times with the dropout layers left active. The predictions produced with the technique are the average of the distribution of outputs produced for each input. Also, a separate model was trained to predict each blood pressure type (systolic/diastolic).
@@
Overall Performance. The best-performing model (XResNet1d50+GNLL on Calib) achieved up to 48% of SBP estimates, and 64% of DBP estimates in the top IEEE category (Grade A), i.e., with prediction errors below 5 mmHg. However, even with this model 41% of SBP estimates and 25% of DBP estimates fell into the lowest category (Grade D). For clinical applications, reducing the fraction of grade D is highly important, or at least identifying patient characteristics that allow to predict whether an unseen patient will belong to grade A or D. This task is referred to as out-of-model-scope detection [49], a task, which is closely related to out-of-distribution detection. Reliable uncertainty estimation is a commonly used approach for the latter and therefore also represents a promising direction for future research. More generally, it might be insightful to assess whether coarser prediction targets, i.e, framing blood pressure prediction as a classification target, could provide clinically meaningful insights, see [50] for an exploration.
@@
Limitations of the VitalDB Dataset. VitalDB is an intraoperative vital signs dataset. Key advantages are that it includes non-cardiac patients, and it is labelled with detailed information on surgical type, surgical approach, anaesthetic type, duration, and device. The dataset covers a wide age range, from neonates to adults. For this study, it was treated as a single dataset, but further disaggregation into subgroups would be essential to identify clinically meaningful features and assess their generalizability in wider healthcare settings. Furthermore, the dataset is not representative of the key application of PPG-based BP monitoring in wearables in the general population, because data were acquired from subjects during surgery rather than in daily life, using clinical pulse oximeters rather than wearable devices.
5.2 Atrial Fibrillation Detection
@@
Relative Performance Comparison. Best performance on the classification case was achieved by modern CNN architectures (XResNet1d, Inception), which performed at least on par with, and in some cases showed advantages over, simpler models (LeNet1d). Feature-based approaches (based on wavelets) showed comparable performance to simple CNNs but were clearly outperformed by large-scale CNNs operating on raw time series input.
@@
Comparison to Literature Results. The original publication describing the DeepBeat dataset [29] reported F1-scores of 0.54 for AF-prediction without pretraining and 0.71 for a multi-task model that jointly predicted AF and signal quality. It is important to stress that these results were achieved on a different, imbalanced split of the dataset and are therefore not directly comparable to those in the present study. The SiamQuality CNN model [46] achieved F1-scores of up to 0.71 on DeepBeat, albeit after self-supervised pretraining. To enable a direct comparison, we trained a XResNet1d50 model on the original split and achieved an F1-score of 0.65, which is clearly superior to the model presented in the original DeepBeat publication. This demonstrates that the presented models reached state-of-the-art performance in comparison to literature benchmarks.
@@
Overall Performance. The best-performing models showed a strong performance in terms of absolute performance values, with F1-scores comparable to or better than reported values in the literature (using the original splits provided by DeepBeat), and sensitivities and specificities of around 0.8. It is worth noting that even though certain commercial AF detection algorithms claim to reach 0.98 sensitivity at >0.99 specificity on internal validation datasets [51], they typically rely on ECG measurements, which serve as the gold standard for AF detection, and severely overestimate the model performance under real-world conditions as the reported performance only applies to certain heart range intervals [52]. In practice, PPG-based AF detection algorithms in consumer wearables are often designed to provide a high positive predictive value by requiring multiple predictions of AF before raising an alert [53, 54] Further work would be required to investigate whether a high positive predictive value could be achieved on real-world data using the models presented in this study.
@@
CIF vs. Raw Time Series Models. The difference in AF detection performance between raw time series models and those based on clinically interpretable features may be influenced by multiple factors, including biases in data acquisition process, artifacts, and even errors in the labelling of recordings into different classes. Deep learning-based raw time series models can use a broad range of information from the signal, enhancing their ability to detect class-specific patterns. However, the drawback is that clinically irrelevant features, such as artifacts and biases, may be incorporated into the training process. In contrast, models that rely on predefined clinically interpretable features, such as pulse irregularity in AF, have a more constrained feature space. While this improves interpretability and robustness in some cases, it may also make these models more vulnerable to mislabelled or noisy data. Initial experiments indicate that incorporating signal quality assessments next to interpretable features can enhance prediction performance. Future work should investigate the classification performance in terms of more fine-grained subgroups, in particular stratified according to signal and labelling quality. Such analysis could reveal if feature-based methods exhibit advantages over raw time series approaches in high-quality subsets, where the underlying assumptions of clinical validity are most clearly satisfied.
@@
Limitations of the DeepBeat Dataset. While the DeepBeat dataset represents one of the few publicly available PPG-based AF prediction datasets which is large enough for training deep learning models, it is very sparsely documented, and no corresponding ECG signals are available, which would allow one to retrospectively assess the annotation quality. For DeepBeat, three independent cohorts were examined: patients admitted for cardioversion before treatment (classified as AF based on patient ID rather than data window),different participants undergoing exercise stress tests and a challenge dataset as the only subset containing both AF and non-AF samples. Furthermore, the data were categorized into low-, medium-, and high-quality segments, all of which were included in this study. However, there was a notable imbalance, with a greater number of high-quality AF-labelled segments compared to non-AF-labelled segments.
@@ 5.3 General Insights and Directions for Future Research
General Recommendations on Model Choices. In this study, modern CNN architectures such as XResNet1d provided the best performance in all three cases. Therefore, our general recommendation is to use these modern CNN architectures for prediction models operating on PPG data. However, in certain cases, it may not be necessary to use such complex models (e.g., see the Calibfree scenario of BP regression). Similarly, for certain tasks, models leveraging image representations achieved competitive results. Nevertheless, in the interest of achieving competitive results on an unknown task, the use of modern CNN architectures based on raw time series representations remains the safest choice.
@@
Future Research Directions. An important restriction of the presented study is the restriction to models that were trained from scratch. A very promising extension lies in the consideration of supervised or self-supervised pretraining, which in the case of DeepBeat lead to an increase from 0.71 to 0.91 in terms of F1-scores for the multi-task model. In particular, this applies to the recently published foundation models for PPG data [55, 56]. A second limitation relates to the restriction of assessing the performance purely based on in-distribution performance, which is known to lead to an overly optimistic assessment of the generalization performance. This urges for dedicated studies on the out-of-distribution generalization performance of the presented models, see [57, 58] for first studies in the context of BP prediction. Finally, the entire investigation focused on quantitative assessment of performance, which neglects other quality dimensions such as interpretability, robustness or uncertainty quantification, all of which merit separate investigations.
@@ 6 Summary and Conclusion
Our investigations across two exemplary regression and classification tasks found that in general modern CNNs (of the ResNet- or Inception-kind) represented the best-performing approaches. The competitiveness of small-scale models and feature-based approaches depends crucially on the definition of the task at hand. While in some scenarios (e.g. regression CalibFree) they can compete with or in some cases even outperform modern CNNs, they may fail to show competitive performance in other cases (e.g. regression Calib, classification).
@@
7 Acknowledgments

### v2

@@ 4.2 AF Detection
T XResNet1d101 0.85 0.66 0.74 0.74 0.53 0.53
XResNet1d50 0.85 0.69 0.75 0.74 0.54 0.54
Inception1d 0.85 0.69 0.74 0.73 0.52 0.52
LeNet1d 0.74 0.55 0.55 0.48 0.34 0.30
AlexNet1d 0.82 0.65 0.68 0.65 0.46 0.45
Minirocket 0.82 0.66 0.69 0.67 0.47 0.47
iTransformer 0.68 0.40 0.47 0.36 0.27 0.17
TimesNet 0.79 0.58 0.61 0.57 0.40 0.38
PPNet 0.85 0.69 0.76 0.75 0.54 0.54
TCN+MLP 0.85 0.67 0.75 0.74 0.54 0.53
F WAVELET + MLP 0.76 0.60 0.58 0.52 0.37 0.33
@@
CWT 0.82 0.69 0.72 0.69 0.50 0.49
I Spectrogram-ResNet-18 0.82 0.68 0.68 0.61 0.47 0.42
Spectrogram-ResNet-50 0.85 0.69 0.72 0.69 0.50 0.49
Table 8: The performance analysis for the classification task on the DeepBeat dataset for three input representations: T for raw time series, F for feature-based, and I for image-based models. The best-performing model is marked in bold and underlined. Models that do not perform statistically significantly worse than the best model are shown in bold.
@@ 5.1 BP Estimation
Relative Performance Comparison. For both regression tasks, the best results were achieved by approaches based on raw time series. Feature-based approaches were not competitive in the Calib scenario, but are to a certain degree competitive in the CalibFree scenario. Here, it is important to note the differences between Calib and CalibFree: CalibFree tests on unseen patients, whereas Calib purposely tests on patients already seen during training, i.e, . models can profit from a certain degree of memorization. Quite naturally, models achieved much better scores in the Calib setting (even though test sets are not entirely comparable). This also impacts the best-forming models in each of the two settings (within the category of models operating on raw time series). Large and complex models (such as XResNet1d and Inception1d) performed best on Calib, as overfitting to specific patients is desirable, whereas smaller models (such as LeNet1d) performed on par or even better than more complex models in the CalibFree scenario, where generalization to unseen patients is key. This observation aligns with the fact that the gap between raw time series and feature-based approaches is larger in the case of Calib, which is quite natural as this task largely profits from the complexity of the model (and all feature-based approaches are less complex than typical models operating on raw time series). Beyond aggregate error metrics, Bland-Altman analysis showed a small degree of systematic bias but large LoA regardless of model, indicating the presence of considerable variability at the sample level even in the best-performing models. This indicates that a smaller MAE does not necessarily imply a reliable prediction at the sample level.
The set of considered raw time series models also included transformer or hybrid CNN-LSTM models. PPNet, in spite of having a fairly modest parameter size ($\sim$125k), was competitive in both Calib and CalibFree cases, coming close to matching the variants of XResNet and emphasizing the potential of light-weight hybrid CNN–LSTM models. Conversely, TimesNet ($\sim$667k parameters) and iTransformer ($\sim$642k parameters) only modestly improved on the baseline and did not reach CNN-level accuracy. Along with the parameter study in Table 10, these indicate that model size does not directly correlate with results: AlexNet (>>40m parameters) and TCNS (>>6m parameters) are among the largest models and yet under-performed, yet the significantly smaller PPNet was more effective.
@@
Comparison to Literature Results. The number of published results on PulseDB is very limited. To the best of our knowledge [16] is the only prior study that reported results on PulseDB, albeit on the full dataset and not just the VitalDB subset, achieving a MAE of 10-11 mmHg for SBP after self-supervised pretraining. We refer to [22] for a recent comparative analysis of different models on different datasets, where models achieved MAEs of between 11 and 19 mmHg for SBP, and between 7 and 11 mmHg for DBP. In comparison, in this study, the lowest MAEs were in the range of approximately 12–13 mmHg and 8 mmHg for SBP and DBP, respectively, on VitalDB CalibFree. This suggests that the presented best-performing models reached state-of-the-art performance.
@@
Impact of Loss Function and Dropout Ensembling. The performance of the XResNet1d50+GNLL in comparison to the XResNet1d50 trained with a regular MAE loss highlights the benefits of incorporating dropout ensembling at evaluation, and using a likelihood-based loss can have on model performance. It suggests that potentially CNN results could be improved by using a different loss function and ensemble-based evaluation procedure. Interestingly, this only applied to the Calib and not the CalibFree scenario. Other works have reported that modelling uncertainties may improve predictive performance on models trained on ECG data [69]. This variant had two parallel output branches with dropout layers (with a constant but low dropout rate, e.g. 4-5% depending on the task/parameter) dispersed throughout the architecture. Each test input was evaluated 100 times with the dropout layers left active. The predictions produced with the technique are the average of the distribution of outputs produced for each input. Also, a separate model was trained to predict each BP type (systolic/diastolic).
@@
Overall Performance. The best-performing model (XResNet1d50+GNLL on Calib) achieved up to 48% of SBP estimates, and 64% of DBP estimates in the top IEEE category (Grade A), i.e. with prediction errors $\leq$ 5 mmHg. However, even with this model, 41% of SBP estimates, and 25% of DBP estimates fell into the lowest category (Grade D). For clinical applications, reducing the fraction of grade D is highly important, or at least identifying patient characteristics that allow to predict whether an unseen patient will belong to grade A or D. This task is referred to as out-of-model-scope detection [24], a task which is closely related to out-of-distribution detection. Reliable uncertainty estimation is a commonly used approach for the latter and therefore also represents a promising direction for future research. More generally, it might be insightful to assess whether coarser prediction targets, i.e. framing BP prediction as a classification target, could provide clinically meaningful insights, see [61] for an exploration.
@@
Limitations of the VitalDB Dataset. VitalDB is an intraoperative vital signs dataset. Key advantages are that it includes non-cardiac patients, and it is labelled with detailed information on surgical type, surgical approach, anaesthetic type, duration, and device. The dataset covers a wide age range, from neonates to adults. For this study, it was treated as a single dataset, but further disaggregation into subgroups would be essential to identify clinically meaningful features and assess their generalizability in wider healthcare settings. Furthermore, the dataset is not representative of the key application of PPG-based BP monitoring in wearables in the general population, because data were acquired from subjects during surgery rather than in daily life, using clinical pulse oximeters rather than wearable devices. It is also worth noting that since the recordings were obtained from anesthetized patients in the controlled surgical settings, the PPG signals contain minimal motion artifacts, and hence, it might result in optimistic performance estimates when compared to real-world wearable scenarios.
5.2 AF Detection
@@
Relative Performance Comparison. Best performance on the classification case was achieved by modern CNN architectures (XResNet1d, Inception), which performed at least on par with, and in some cases showed advantages over, simpler models (LeNet1d). Feature-based approaches (based on wavelets) showed comparable performance to simple CNNs but were clearly outperformed by large-scale CNNs operating on raw time series input. Furthermore, it is worth noting that although PPNet’s performance was very close to the top-performing CNNs, it still fell short, whereas both iTransformer and TimesNet were distinctly inferior, highlighting the relative effectiveness of the convolutional architectures on this task. Notably, the STFT-based Spectrogram-ResNet achieved performance comparable to the top-performing raw time-series CNNs in terms of AUC, F1-score, and thresholded sensitivity and specificity, indicating that image-based representations can be competitive for AF detection. Furthermore, bootstrap analyses showed that several top-performing CNN and image-based models were statistically indistinguishable in terms of classification accuracy, suggesting that increasing architectural complexity yields diminishing returns for this data.
@@
Comparison to Literature Results. The original publication describing the DeepBeat dataset [67] reported F1-scores of 0.54 for AF-prediction without pretraining and 0.71 for a multi-task model that jointly predicted AF and signal quality. It is important to stress that these results were achieved on a different, imbalanced split of the dataset and are therefore not directly comparable to those in the present study. The SiamQuality CNN model [16] achieved F1-scores of up to 0.71 on DeepBeat, albeit after self-supervised pretraining. To enable a direct comparison, we train an XResNet1d50 model on the original split and achieved an F1-score of 0.69, which is clearly superior to the model presented in the original DeepBeat publication. This demonstrates that the presented models reached state-of-the-art performance in comparison to literature benchmarks.
@@
Overall Performance. The best-performing models showed a strong performance in terms of absolute performance values, with F1-scores comparable to or better than reported values in the literature (using the original splits provided by DeepBeat), and sensitivities and specificities of around 0.8. It is worth noting that even though certain commercial AF detection algorithms claim to reach 0.98 sensitivity at >> 0.99 specificity on internal validation datasets [62], they typically rely on ECG measurements, which serve as the gold standard for AF detection, and severely overestimate the model performance under real-world conditions as the reported performance only applies to certain heart range intervals [27]. In practice, PPG-based AF detection algorithms in consumer wearables are often designed to provide a high positive predictive value by requiring multiple predictions of AF before raising an alert [42, 55]. Further work would be required to investigate whether a high positive predictive value could be achieved on real-world data using the models presented in this study.
@@
CIF vs. Raw Time Series Models. The difference in AF detection performance between raw time series models and those based on clinically interpretable features may be influenced by multiple factors, including biases in the data acquisition process, artifacts, and even errors in the labelling of recordings into different classes. Deep learning-based raw time series models can use a broad range of information from the signal, enhancing their ability to detect class-specific patterns. However, the drawback is that clinically irrelevant features, such as artifacts and biases, may be incorporated into the training process. In contrast, models that rely on predefined clinically interpretable features, such as pulse irregularity in AF, have a more constrained feature space. While this improves interpretability and robustness in some cases, it may also make these models more vulnerable to mislabelled or noisy data. Initial experiments indicate that incorporating signal quality assessments next to interpretable features can enhance prediction performance. Future work should investigate the classification performance in terms of more fine-grained subgroups, in particular stratified according to signal and labelling quality. Such analysis could reveal if feature-based methods exhibit advantages over raw time series approaches in high-quality subsets, where the underlying assumptions of clinical validity are most clearly satisfied.
@@
Limitations of the DeepBeat Dataset. While the DeepBeat dataset represents one of the few publicly available PPG-based AF prediction dataset that is large enough for training deep learning models, it is very sparsely documented, and no corresponding ECG signals are available, which would allow one to retrospectively assess the annotation quality. For DeepBeat, three independent cohorts were examined: patients admitted for cardioversion before treatment (classified as AF based on patient ID rather than data window), different participants undergoing exercise stress tests and a challenge dataset as the only subset containing both AF and non-AF samples. Furthermore, the data were categorized into low-, medium-, and high-quality segments, all of which were included in this study. However, there was a notable imbalance, with a greater number of high-quality AF-labelled segments compared to non-AF-labelled segments.
@@ 5.3 General Insights and Directions for Future Research
General Recommendations on Model Choices. In this study, modern CNN architectures such as XResNet1d provided the best performance in all three cases, as confirmed by both aggregate performance metrics and agreement- and bootstrap-based analyses. Therefore, our general recommendation is to use these modern CNN architectures for prediction models operating on PPG data. However, in certain cases, it may not be necessary to use such complex models (e.g. see the Calibfree scenario of BP regression). Similarly, for certain tasks, models leveraging image representations achieved competitive results. Nevertheless, in the interest of achieving competitive results on an unknown task, the use of modern CNN architectures based on raw time series representations remains the safest choice. In spite of promising results in other applications, more complicated hybrid CNN-RNN or transformer-based models did not show any improvements over CNN-based baselines. This might not be a finding that generalizes to other datasets, but potentially just a reflection of the fact that the considered models for the respective datasets/tasks are already operating in a regime where they are largely data-constrained and already close to the optimal performance achievable on the respective datasets.
Model Interpretability. Beyond quantitative performance metrics, model interpretability represents a critical factor determining clinical adoption of AI systems. This interpretability challenge manifests differently across modelling approaches. Feature-based methods offer inherent interpretability advantages through clinically meaningful inputs such as pulse morphology metrics or rhythm irregularity characteristics. These inputs provide immediate physiological interpretation that clinicians can readily understand. However, when complex prediction algorithms operate on these interpretable features, the overall model often becomes opaque, negating the initial interpretability benefit.
The explainable AI (XAI) community has developed techniques to address this opacity across different input representations. Post-hoc XAI methods, applicable to already-trained models, commonly generate heatmaps that highlight input regions contributing positively or negatively to model decisions [30]. These approaches have shown promise when applied to physiological time series analysis [70]. Despite advances in XAI techniques, deep learning models operating on raw data remain fundamentally more challenging to interpret than models built on clinically meaningful features. This interpretability gap persists even with sophisticated explanation methods, creating a fundamental trade-off between model sophistication and clinical transparency.
Robustness. There is a second important quality dimension that is not captured explicitly by our analysis: robustness to input noise or artefacts. Note that wrist-worn PPG sensors, as used in DeepBeat, are more prone to motion artefacts and provide a lower signal-to-noise ratio relative to clinical-grade systems. But also within a single dataset, signal quality varies widely. In free-living conditions, motion artefacts often dominate wrist PPG recordings and represent one of the primary reasons why many academic PPG-based models fail to translate into reliable real-world devices. Future work should evaluate the robustness, for example, by considering predictive performance stratified according to signal quality. While datasets such as DeepBeat or MIMIC would, in principle, allow such analyses, the present study focuses on controlled, standardized evaluation settings to enable like-for-like comparison of model architectures, and therefore does not include an explicit quantitative robustness assessment. A second robustness limitation relates to the restriction of assessing the performance purely based on in-distribution performance, which is known to lead to an overly optimistic assessment of the generalization performance. This urges for dedicated studies on the out-of-distribution generalization performance of the presented models, see [74, 48] for the first studies in the context of BP prediction.
@@
(Self-supervised) Pretraining. An important restriction of the presented study is the restriction to models that were trained from scratch. A very promising extension lies in the consideration of supervised or self-supervised pretraining, which, in the case of DeepBeat, leads to an increase from 0.71 to 0.91 in terms of F1-scores for the multi-task model. In particular, this applies to the recently published foundation models for PPG data [16, 57].
@@ 6 Summary and Conclusion
Our investigations across two exemplary regression and classification tasks found that in general modern CNNs (of the ResNet- or Inception-kind) represented the best-performing approaches. The competitiveness of small-scale models and feature-based approaches depends crucially on the definition of the task at hand. While in some scenarios (e.g. regression CalibFree), they can compete with or in some cases even outperform modern CNNs, they may fail to show competitive performance in other cases (e.g. regression Calib, classification).
Code availability
@@
Generative AI statement
No generative AI was used for text generation in this manuscript.
Competing interests
The authors have declared no competing interests.
Ethics statement
This study used publicly available and deidentified datasets. No direct patient interaction or intervention was involved. As these datasets are released under established data use agreements and have been ethically approved for secondary research, no additional ethics approval was sought.
Acknowledgments